OpenStudy (anonymous):

What’s the equation of a line that passes through points (0, -1) and (2, 3)? can sum1 plz show me how..with formulass

OpenStudy (anonymous):

y = mx+b slope: (y2-y1)/(x2-x1) (x1,y1) = (0,-1) (x2,y2) = (2,3) slope: (3-(-1))/((2-0) = 4/2 = 2 we know the y intercept is when x = 0. we have that point (0,-1) so b is -1. y = 2x+(-1) so y = 2x-1